# Flow-based Domain Randomization for Learning and Sequencing Robotic Skills

### v1

@@ Abstract
Domain randomization in reinforcement learning is an established technique for increasing the robustness of control policies trained in simulation. By randomizing environment properties during training, the learned policy can become robust to uncertainties along the randomized dimensions. While the environment distribution is typically specified by hand, in this paper we investigate automatically discovering a sampling distribution via entropy-regularized reward maximization of a normalizing-flow–based neural sampling distribution. We show that this architecture is more flexible and provides greater robustness than existing approaches that learn simpler, parameterized sampling distributions, as demonstrated in six simulated and one real-world robotics domain. Lastly, we explore how these learned sampling distributions—combined with a privileged value function—can be used for out-of-distribution detection in an uncertainty-aware multi-step manipulation planner.
@@ 1 Introduction
Reinforcement learning (RL) has proven to be a useful tool in robotics for learning control or action policies for tasks and systems which are highly variable and/or analytically intractable (Luo & Li, 2021; Zhu et al., 2020; Schoettler et al., 2020). However, RL approaches can be inefficient, involving slow, minimally parallelized, and potentially unsafe data-gathering processes when performed in real environments (Kober et al., 2013). Learning in simulation eliminates some of these problems, but introduces new issues in the form of discrepancies between the training and real-world environments (Valassakis et al., 2020).
@@
An ideal sampling distribution enables the policy to focus training on areas of the distribution that can feasibly be solved in order to maximize the overall success rate of the trained policy while not wasting time on unsolvable regions of the domain. Automating updates to parameter distributions during the training process can remove the need for heuristic tuning and iterative experimentation (Mozifian et al., 2019; OpenAI et al., 2019; Tiboni et al., 2024). In this paper, we present GoFlow, a novel approach for learned domain randomization that combines actor-critic reinforcement learning architectures (Schulman et al., 2017; Haarnoja et al., 2018) with a neural sampling distribution to learn robust policies that generalize to real-world settings. By maximizing the diversity of parameters during sampling, we actively discover environments that are challenging for the current policy but still solvable given enough training.
@@
Lastly, we extend this classical gear insertion task to the setting of multi-step decision making under uncertainty and partial observability. As shown in this paper and elsewhere (Tiboni et al., 2024; Mozifian et al., 2019; OpenAI et al., 2019), policies trained in simulation have an upper bound on the environmental uncertainties that they can be conformant to. For example, a visionless robot executing an insertion policy can only tolerate so much in-hand pose error. However, estimates of this uncertainty can be used to inform high level control decisions, e.g., looking closer at objects to get more accurate pose estimates or tracking objects in the hand to detect slippage. By integrating a probabilistic pose estimation model, we can use the sampling distributions learned with GoFlow as an out-of-distribution detector to determine whether the policy is expected to succeed under its current belief about the world state. If the robot has insufficient information, it can act to deliberately seek the needed information using a simple belief-space planning algorithm. For example, an in-hand camera can be used to gather a higher resolution image for more accurate pose estimation if necessary.
Our contributions are as follows: We introduce GoFlow, a novel domain randomization method that combines actor-critic reinforcement learning with a learned neural sampling distribution. We show that GoFlow outperforms fixed and other learning-based solutions to domain randomization on a suite of simulated environments. We demonstrate the efficacy of GoFlow in a real-world contact-rich manipulation task—gear insertion—and extend it to multi-step decision-making under uncertainty. By integrating a probabilistic pose estimation model, we enable the robot to actively gather additional information when needed, enhancing performance in partially observable settings.
@@ 2 Related Work
Recent developments in reinforcement learning have proven that policies trained in simulation can be effectively translated to real-world robots for contact-rich assembly tasks (Zhang et al., 2024; Tang et al., 2023a; Noseworthy et al., 2024; Jin et al., 2023). One key innovation that has contributed to the development of robust policies is domain randomization (Chen et al., 2021; Peng et al., 2017), wherein environment parameters are sampled from a distribution during training such that the learned policy can be robust to environmental uncertainty on deployment.
Some previously explored learning strategies include minimization of divergence with a target sampling distribution using multivariate Gaussians (Mozifian et al., 2019), maximization of entropy using independent beta distributions (Tiboni et al., 2024), and progressive expansion of a uniform sampling distribution via boundary sampling (OpenAI et al., 2019). Here, we propose a novel learned domain randomization technique using normalizing flows (Rezende & Mohamed, 2015) as a neural sampling distribution, thus increasing flexibility and expressivity.
In addition to learning robust policies, such sampling distributions can be used as indicators of the world states under which the policy is expected to succeed. Some previous works have combined domain randomization with information gathering via system identification (Ramos et al., 2019; Sagawa & Hino, 2024). In this work, we similarly make use of our learned sampling distribution as an out-of-distribution detector in the context of a multi-step planning system.
@@ 5.1 Domains
Second, we quantitatively compare GoFlow to existing baselines including Cartpole, Ant, Quadcopter, and Quadruped in the IsaacLab suite of environments (Mittal et al., 2023). We randomize over parameters such as link masses, joint frictions, and material properties.
@@ 5.2 Baselines
In our domain randomization experiments, we compare to a number of standard RL baselines and learning-based approaches from the literature. In our quantitative experiments, success is measured by the sampled environment passing a certain performance threshold $J_{T}$ that was selected for each environment. All baselines are trained with an identical neural architecture and PPO implementation. The success thresholds along with other hyperparameters are in Appendix A.5.
@@
We measure task performance in terms of coverage, which is defined as the proportion of the total sampling distribution for which the policy receives higher than $J_{t}$ reward. We compare coverage across environments sampled from a uniform testing distribution within the environment bounds. Our findings in Figure 3 show that GoFlow matches or outperforms baselines across all domains. We find that our method performs particularly well in comparison to other learned baselines when the simpler or more feasible regions of the domain are off-center, irregularly shaped, and have inter-parameter dependencies such as those seen in Figure 2. We intentionally chose large intervals with low coverage to demonstrate this capability, and we perform a full study of how baselines degrade significantly with increased parameter ranges while GoFlow degrades more gracefully (see Appendix A.6). Additionally, coverage should not be mistaken for success-rate, as these policies can be integrated into a planning framework that avoids the use of infeasible actions or gathers information to increase coverage as discussed in the following sections.
@@ 6.1 Belief-space planning background
Figure 4: A multi-step manipulation plan using probabilistic pose estimation to estimate and update beliefs over time. The three rows show the robot state $s_{t}$, the observation $o_{t}$, and the robot belief $b_{t}$ at each timestep. The red dotted line in the belief indicates the marginal entropy thresholds for the x, y, and yaw dimensions as determined by the learned normalizing flow. For full visualizations of the belief posteriors, flow distributions, and value maps, see Figure 13 in the appendix.
@@ 6.3 Updating beliefs
Figure 5: A visual example of the precondition computation described in Section 6.2 for a single step in the gear assembly process. The normalizing flow is thresholded to indicate the in-distribution regions of the learned value function. The thresholded sampling distribution is further thresholded by the value function to get the belief-space precondition. Comparing the precondition to the beliefs, we can see that the belief is not sufficiently contained within the precondition at $t=0$, but passes the success threshold $\eta$ at after closer inspection $t=4$.
@@ 6.4 A simple belief-space planner
An example plan can be seen in Figure 4. The goal is to assemble the gear box by inserting all three gears (yellow, pink, and blue) into the shafts on the gear plate. Each gear insertion is associated with a separate policy for each color trained with GoFlow. In addition to the trained policies, the robot is given access to an object-parameterized inspection action which has no preconditions and whose effects are a reduced-variance pose estimate attained by moving the camera closer to the object. The robot is initially uncertain of the x, y, and yaw components of the 6-dof pose based on probabilistic pose estimates. Despite this uncertainty, the robot is confident enough in the pose of the largest and closest yellow gear to pick it up and insert it. In contrast, the blue and pink gears require further inspection to get a better pose estimate. Closer inspection reduces uncertainty along the x and y axis, but reveals no additional information about yaw dimension due to rotational symmetry. Despite an unknown yaw dimension, the robot is confident in the insertion because the flow $p_{\phi}$ indicates that success is invariant to the yaw dimension. For visualizations of the beliefs and flows at each step, see Appendix A.4.
@@ 7 Conclusion and discussion
In this paper, we introduced GoFlow, a novel approach to domain randomization that uses normalizing flows to dynamically adjust the sampling distribution during reinforcement learning. By combining actor-critic reinforcement learning with a learned neural sampling distribution, we enabled more flexible and expressive parameterization of environmental variables, leading to better generalization in complex tasks like contact-rich assembly. Our experiments demonstrated that GoFlow outperforms traditional fixed and learning-based domain randomization techniques across a variety of simulated environments, particularly in scenarios where the domain has irregular dependencies between parameters. The method also showed promise in real-world robotic tasks including contact-rich assembly.
@@ Appendix A Appendix
A.1 Reproducibility Statement
In our supplementary materials, we provide a minimal code implementation of our approach along with all the baseline implementations and environments discussed in this paper. If accepted, we will publish a full public release of the code on Github.
@@ A.3 Domain Randomization Parameters
Below we describe the randomization ranges and parameter names for each environment. We also provide the reward success threshold ($J_{T}$) and cut the max duration of some environments in order to speed up training ($t_{max}$). Lastly, we slightly modified the Quadruped environment to only take a fixed forward command rather than the goal-conditioned policy learned by default. Other than those changes, the first four simulated environments official IsaacLab implementation.
@@ A.5.4 ADR
A.6 Coverage vs Range Experiments
Figure 12: Coverage vs range experiment results
A.7 Real-world experiments
Method Success Rate
FullDR 6/10
NoDR 3/10
DORAEMON 5/10
LSDR 5/10
ADR 5/10
GoFlow 9/10
@@ A.7 Real-world experiments
FullDR .017$\pm$.012 .003$\pm$.002 .003$\pm$.004 .$0$200$\pm$.198 .001$\pm$.000
NoDR .$0$$0$$0$062$\pm$.048 .$0$010$\pm$.000
DORAEMON .$0$$0$$0$023$\pm$.022 .124$\pm$.248 .020$\pm$.018
LSDR .110$\pm$.006 .$0$002$\pm$.002 .011$\pm$.002 .278$\pm$.230 .003$\pm$.000
ADR .$0$$0$$0$114$\pm$.026 .$0$072$\pm$.044
GoFlow .109$\pm$.080 .072$\pm$.038 .019$\pm$.010 .178$\pm$.220 .321 $\pm$ .068 .104$\pm$.028
Table 2: Mean and standard deviation of coverage, with all statistically significant entries bolded.

### v2

@@ Abstract
Domain randomization in reinforcement learning is an established technique for increasing the robustness of control policies trained in simulation. By randomizing environment properties during training, the learned policy can become robust to uncertainties along the randomized dimensions. While the environment distribution is typically specified by hand, in this paper we investigate automatically discovering a sampling distribution via entropy-regularized reward maximization of a normalizing-flow–based neural sampling distribution. We show that this architecture is more flexible and provides greater robustness than existing approaches that learn simpler, parameterized sampling distributions, as demonstrated in six simulated and one real-world robotics domain. Lastly, we explore how these learned sampling distributions, along with a privileged value function, can be used for out-of-distribution detection in an uncertainty-aware multi-step manipulation planner.
@@ 1 Introduction
Reinforcement learning (RL) is a powerful tool in robotics because it can be used to learn effective control policies for systems with highly complex dynamics that are difficult to model analytically. Unlike traditional control methods, which rely on precise mathematical models, RL learns directly from simulated or real-world experience (Luo & Li, 2021; Zhu et al., 2020; Schoettler et al., 2020).
However, RL approaches can be inefficient, involving slow, minimally parallelized, and potentially unsafe data-gathering processes when performed in real environments (Kober et al., 2013). Learning in simulation eliminates some of these problems, but introduces new issues in the form of discrepancies between the training and real-world environments (Valassakis et al., 2020).
@@
An alternative strategy is to learn an environment distribution during training with the aim of finding the broadest possible training distribution that can feasibly be solved in order to maximize the chances of transferring to an unknown target environment. Automating updates to parameter distributions during the training process can remove the need for heuristic tuning and iterative experimentation (Mozifian et al., 2019; OpenAI et al., 2019; Tiboni et al., 2024). In this paper, we present GoFlow, a novel approach for learned domain randomization that combines actor-critic reinforcement learning architectures (Schulman et al., 2017; Haarnoja et al., 2018) with a neural sampling distribution to learn robust policies that generalize to real-world settings. By maximizing the diversity of parameters during sampling, we actively discover environments that are challenging for the current policy but still solvable given enough training.
@@
Lastly, we extend this classical gear insertion task to the setting of multi-step decision making under uncertainty and partial observability. As shown in this paper and elsewhere (Tiboni et al., 2024; Mozifian et al., 2019; OpenAI et al., 2019), policies trained in simulation have an upper bound on the environmental uncertainties that they can be conformant to. For example, a visionless robot executing an insertion policy can only tolerate so much in-hand pose error. However, estimates of this uncertainty can be used to inform high level control decisions through task-oriented information gathering (Liang et al., 2020; Curtis et al., 2022). For example, looking closer at objects can result in more accurate pose estimates or tracking objects in the hand to detect slippage. By integrating a probabilistic pose estimation model, we can use the sampling distributions learned with GoFlow as an out-of-distribution detector to determine whether the policy is expected to succeed under its current belief about the world state. If the robot has insufficient information, it can act to deliberately seek the needed information using a simple belief-space planning algorithm.
Our contributions are as follows: We introduce GoFlow, a novel domain randomization method that combines actor-critic reinforcement learning with a learned neural sampling distribution. We show that GoFlow achieves higher domain coverage than fixed and other learning-based solutions to domain randomization on a suite of simulated environments. We demonstrate the efficacy of GoFlow in a real-world contact-rich manipulation task—gear insertion—and extend it to multi-step decision-making under uncertainty. By integrating a probabilistic pose estimation model, we enable the robot to actively gather additional information when needed, enhancing performance in partially observable settings.
@@ 2 Related Work
Recent developments in reinforcement learning have demonstrated that policies trained in simulation can effectively transfer to real-world robots, even for contact-rich robotic tasks (Zhang et al., 2024; Tang et al., 2023a; Noseworthy et al., 2024; Jin et al., 2023). A key innovation enabling this transferability is domain randomization (Chen et al., 2021; Peng et al., 2017), where environment parameters are sampled from predefined distributions during training, enabling learned policies to generalize to environmental uncertainties upon deployment.
Traditionally, domain randomization requires manually defining these sampling distributions, which can be labor-intensive. To address this, recent work has explored methods to automatically learn these distributions, aiming for maximal generalization with minimal manual effort. One notable approach involves constructing adversarial distributions that challenge the current policy, ensuring coverage of the environment parameter space (Mehta et al., 2020; Wang et al., 2025). While adversarial methods can outperform uniform randomization, they typically assume all environments sampled are solvable, limiting their effectiveness in highly uncertain scenarios.
An alternative line of research adopts a curriculum-based approach, progressively expanding the complexity of the training distribution while maintaining policy success. Specific curricular methods include minimizing divergence from a target distribution using multivariate Gaussians (Mozifian et al., 2019), maximizing entropy through independent beta distributions (Tiboni et al., 2024), and incrementally expanding uniform sampling distributions via boundary sampling (OpenAI et al., 2019). In this work, we propose a novel learned domain randomization technique leveraging normalizing flows (Rezende & Mohamed, 2015) as neural sampling distributions. This approach offers increased flexibility and expressivity over existing parametric methods.
Beyond training robust policies in simulation, learned sampling distributions can be tied to the real-world environmental conditions under which policies are likely to succeed. Previous works have integrated domain randomization with real-world interactions for more informed training distributions (Ramos et al., 2019; Sagawa & Hino, 2024; Ajay et al., 2023; Muratore et al., 2020) or to find the maximally effective real-world strategy  (Yu et al., 2018; Ren et al., 2023). However, these methods often necessitate expensive policy retraining or data-intensive evolutionary search based on real-world feedback, posing challenges for real-time applications. Instead, we utilize our learned sampling distribution as an out-of-distribution detector within a multi-step planning framework, enabling fast and data-efficient information gathering in the real world.
@@ 5.1 Domains
Second, we quantitatively compare GoFlow to existing baselines including Cartpole, Ant, Quadcopter, Quadruped, and Humanoid in the IsaacLab suite of environments (Mittal et al., 2023). We randomize over parameters such as link masses, joint frictions, and material properties.
@@ 5.2 Baselines
In our domain randomization experiments, we compare to a number of standard RL baselines and learning-based approaches from the literature. In our quantitative experiments, success is determined by the sampled environment passing a certain performance threshold $J_{T}$ that was selected for each environment. We measure task performance in terms of coverage, which is defined as the proportion of the total sampling distribution for which the policy receives higher than $J_{t}$ reward. Coverage is estimated via policy rollouts on 4096 uniformly selected environment initialization samples from the total range of parameters. All baselines are trained with an identical neural architecture and PPO implementation. The success thresholds along with other hyperparameters are in Appendix A.5.
@@
We compare coverage across environments sampled from a uniform testing distribution within the environment bounds. Our findings in Figure 3 show that GoFlow matches or outperforms baselines across all domains. We find that our method performs particularly well in comparison to other learned baselines when the simpler or more feasible regions of the domain are off-center, irregularly shaped, and have inter-parameter dependencies such as those seen in Figure 2. We intentionally chose large intervals with low coverage to demonstrate this capability, and we perform a full study of how baselines degrade significantly with increased parameter ranges while GoFlow degrades more gracefully (see Appendix A.6). Additionally, coverage should not be mistaken for success-rate, as these policies can be integrated into a planning framework that avoids the use of infeasible actions or gathers information to increase coverage as discussed in the following sections.
@@
Figure 4: A multi-step manipulation plan using probabilistic pose estimation to estimate and update beliefs over time. The three rows show the robot state $s_{t}$, the observation $o_{t}$, and the robot belief $b_{t}$ at each timestep. The red dotted line in the belief indicates the marginal entropy thresholds for the x, y, and yaw (rotation around z) dimensions as determined by the learned normalizing flow. A belief with entropy surpassing the threshold line indicates the policy will likely fail. For full visualizations of the belief posteriors, flow distributions, and value maps, see Figure 14.
@@ 6.2 Computing preconditions
Figure 5: A visual example of the precondition computation described in Section 6.2 for the gear assembly plan shown in Figure 4. The two rows show two different projections of the 3D sampling space (x position vs y position in the top row and y position vs yaw rotation in the bottom row). We apply a threshold $\epsilon$ to the sampling distribution to remove low-probability regions (column 1). Additionally, we filter the value function by retaining only the regions where the expected value exceeds a predetermined threshold $\eta$ (column 2). The intersection of these two regions defines the belief-space precondition, indicating where the policy is likely to succeed (column 3). Comparing the precondition to the beliefs, we can see that the belief is not sufficiently contained within the precondition at $t=0$ (column 4), but passes the success threshold $\eta$ at after closer inspection at $t=4$ (column 5).
@@ 6.4 A simple belief-space planner
An example plan can be seen in Figure 4. The goal is to assemble the gear box by inserting all three gears (yellow, pink, and blue) into the shafts on the gear plate. Each gear insertion is associated with a separate policy for each color trained with GoFlow. In addition to the trained policies, the robot is given access to an object-parameterized inspection action which has no preconditions and whose effects are a reduced-variance pose estimate attained by moving the camera closer to the object. The robot is initially uncertain of the x, y, and yaw components of the 6-dof pose based on probabilistic pose estimates. Despite this uncertainty, the robot is confident enough in the pose of the largest and closest yellow gear to pick it up and insert it. In contrast, the blue and pink gears require further inspection to get a better pose estimate. Closer inspection reduces uncertainty along the x and y axis, but reveals no additional information about yaw dimension due to rotational symmetry. Despite an unknown yaw dimension, the robot is confident in the insertion because the flow $p_{\phi}$ indicates that success is invariant to the yaw dimension. This is due to the fact that success in the insertion task is defined by the distance between the bottom center of the gear and the base of the gear shaft, which is independent of gear rotation. For visualizations of the beliefs and flows at each step, see Appendix A.4.
@@ 7 Conclusion and discussion
In this paper, we introduced GoFlow, a novel approach to domain randomization that uses normalizing flows to dynamically adjust the sampling distribution during reinforcement learning. By combining actor-critic reinforcement learning with a learned neural sampling distribution, we enabled more flexible and expressive parameterization of environmental variables, leading to better generalization in complex tasks like contact-rich assembly. Our experiments demonstrated that GoFlow achieves higher coverage than traditional fixed and learning-based domain randomization techniques across a variety of simulated environments, particularly in scenarios where the domain has irregular dependencies between parameters. The method also showed promise in real-world robotic tasks including contact-rich assembly.
@@ Appendix A Appendix
A.1 Code Release
The codebase for the project can be found here.
@@ A.3 Domain Randomization Parameters
Below we describe the randomization ranges and parameter names for each environment. We also provide the reward success threshold ($J_{T}$) and cut the max duration of some environments in order to speed up training ($t_{max}$). $J_{t}$ was chosen to be below the optimal performance under no environment randomization. We verified that the trained policy still exhibited qualitatively successful performance at the target reward threshold. Lastly, we slightly modified the Quadruped environment to only take a fixed forward command rather than the goal-conditioned policy learned by default. Other than those changes, the first four simulated environments official IsaacLab implementation.
@@ A.5.1 GoFlow
A.5.2 LSDR
A.5.3 DORAEMON
A.5.4 ADR
A.6 Coverage vs Range Experiments
A.7 Coverage vs. Threshold Experiments
In this experiment, we assess the sensitivity of coverage to the performance threshold $J_{t}$ for a single training seed. We experiment with threshold multipliers between 00 and $1.2\cdot{J_{t}}$. The results of this experiment can be seen in Figure 13. These results demonstrate that although coverage is highly dependent on $J_{t}$, GoFlow largely dominates the coverage distribution for wide ranges of $J_{t}$.
A.8 Real-world experiments
A.9 Statistical Tests
@@
Figure 12: Coverage vs range experiment results discussed in Section A.6
Figure 13: Coverage vs threshold experiment results discussed in Section A.7
Method FullDR NoDR DORAEMON LSDR ADR GoFlow
Success Rate 6/10 3/10 5/10 5/10 5/10 9/10
@@
FullDR 0.060$\pm$0.021 0.014$\pm$0.006 0.000$\pm$0.000 0.012$\pm$0.007 0.195$\pm$0.044 0.000$\pm$0.000
NoDR 0.005$\pm$0.001 0.000$\pm$0.000 0.000$\pm$0.000 0.086$\pm$0.008 0.001$\pm$0.001 0.008$\pm$0.003
DORAEMON 0.015$\pm$0.005 0.000$\pm$0.000 0.000$\pm$0.000 0.029$\pm$0.008 0.000$\pm$0.000 0.012$\pm$0.004
LSDR 0.103$\pm$0.013 0.000$\pm$0.000 0.007$\pm$0.003 0.009$\pm$0.008 0.176$\pm$0.082 0.000$\pm$0.000
ADR 0.006$\pm$0.001 0.000$\pm$0.000 0.000$\pm$0.000 0.070$\pm$0.006 0.000$\pm$0.000 0.008$\pm$0.004
GOFLOW 0.138$\pm$0.005 0.056$\pm$0.033 0.010$\pm$0.006 0.275$\pm$0.019 0.274$\pm$0.076 0.035$\pm$0.019
Table 2: Mean and standard error (SDE) of the final reward value, with entries in bold indicating the best-performing method or methods not statistically significantly worse than the best (one-tailed z test, $\alpha=0.05$).
Cartpole Ant Quadcopter Quadruped Humanoid Gears
FullDR -920.132 362.454 -2.286 0.108 657.414 8.627
NoDR -949.925 133.331 -3.807 -4.062 180.734 12.697
DORAEMON -930.417 86.742 -1.767 -2.472 235.516 11.378
LSDR -867.347 367.612 -1.870 -0.052 532.531 9.289
ADR -953.498 64.815 -3.183 -2.926 246.297 9.671
GOFLOW -820.510 198.302 -1.000 -2.814 376.652 14.209
Table 3: CVaR computed as the mean of the final rewards falling below the 10% percentile (VaR).